# The dynamical and thermodynamic effects of turbulence for the cosmic baryonic fluid

### v1

@@ 4 Conclusions and Discussions
Turbulence in the cosmic baryonic fluid can exert both dynamical and thermodynamic influences on cosmic baryons. For the dynamical aspect of turbulence, we find that in both the low- and intermediate-density regions, the dynamical effects are negative and dominated by turbulence, that is $\overline{Q}_{\rm tot}\sim\overline{Q}_{\rm turb}<0$. These negative dynamical effects suppress outflow in low-density regions but promote inflow in intermediate-density regions, which causes the gas to be detained in these regions. Conversely, in high-density regions, the dynamical effects are positive and dominated by thermal pressure, that is $\overline{Q}_{\rm tot}\sim\overline{Q}_{\rm th}>0$, which will prevent the inflow of the gas. These dynamical effects result in high baryon fractions in low- and intermediate-density regions, but low baryon fractions in high-density regions. Furthermore, the cumulative mass fraction reveals that both the WHIM and warm gas are more extensively distributed in space.
@@
For the thermodynamic aspect of turbulence, turbulent energy is converted into thermal energy in two primary ways. Firstly, turbulent kinetic energy cascades from large-scale to small-scale eddies, dissipating into heat when the scale falls below the Kolmogorov length, which is suitable for Kolmogorov turbulence. Secondly, if the power index of turbulent energy spectra is steeper than the Kolmogorov index, kinetic energy can be directly dissipated into thermal energy via shock heating within the inertial range. Consequently, the energy dissipation rate for turbulence with a power index steeper than $-5/3$ exceeds that of Kolmogorov turbulence (Wang & He, 2024a). It can be seen, from $z=1$ to $z=0$, that both the mean thermal energy density $\overline{\varepsilon}_{\rm th}$ and the mean turbulent energy density $\overline{\varepsilon}_{\rm turb}$ remain nearly constant, and the shapes and amplitudes of the environment-dependent wavelet energy spectra are essentially unchanged. These findings indicate that the injection and dissipation of turbulent energy have essentially reached a balance, and the cosmic fluid is in a state of fully-developed turbulence within this redshift range.
@@
Due to turbulent heating, we see that as redshift decreases, an increasing amount of warm gas is heated and transitions into the WHIM, and some even into hot gas. Consequently, the mass fraction of warm gas gradually decreases, while that of the WHIM increases. The WHIM is distributed in nearly all regions, with the majority concentrated in intermediate-density regions.
Additionally, a related issue is the overcooling problem (Voit, 2005). If the gas cools too much, it could lead to an overabundance of stars and galaxies forming at early times, which is inconsistent with observational data. Thus, feedback from SN and AGN is proposed to resolve the problem. However, as addressed in Silk & Nusser (2010), black holes cannot supply sufficient momentum in radiation to expel the gas by pressure alone. We propose that, in addition to SN feedback and AGN activities, turbulent heating should also be a potential heating mechanism to overcome the overcooling problem.
In summary, through dynamical and thermodynamic studies of the cosmic baryonic fluid, we have provided a physical explanation for the missing baryon problem.
Acknowledgements
@@
We acknowledge the use of the data from IllustrisTNG simulation for this work. We also acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).

### v3

@@ 4 Conclusions and discussions
Turbulence in the cosmic baryonic fluid can exert both dynamical and thermodynamic influences on cosmic baryons. According to the dark matter density, we divided the simulation space into the following regions: (1) low-density, (2) intermediate-density, (3) high-density, and (4) extremely high-density. This roughly correspond to (1) voids and under-dense regions, (2) filaments, sheets, and the outskirts of clusters in the cosmic web, (3) virialized structures such as galaxy clusters, and (4) galaxies in clusters. Similar to Bregman (2007), we also defined four components of baryonic matter: (1) hot gas, (2) the WHIM, (3) warm gas, and (4) the cold component. In this work, using the IllustrisTNG50-1 data, we systematically investigated the dynamical and thermodynamic effects of turbulence in cosmic baryonic fluids across different density regions. Our main findings and conclusions are as follows:
  1. 1.
@@
    For the dynamical aspect, we find that in both low- and intermediate-density regions, the dynamical effects are negative and dominated by turbulence, that is, $\overline{Q}_{\rm tot}\sim\overline{Q}_{\rm turb}<0$. These negative dynamical effects suppress expansion in low-density regions but promote convergence in intermediate-density regions, which causes the gas to be detained in these regions. Conversely, in high-density regions, the dynamical effects are positive and dominated by thermal pressure, that is, $\overline{Q}_{\rm tot}\sim\overline{Q}_{\rm th}>0$, which prevents the convergence of the gas. In all cases, the dynamical effects of both turbulence and shocks are negative. In low-density regions the mean dynamical effects of shocks are marginally smaller than, yet comparable to, those of turbulence. However, in intermediate-density regions, the effects of turbulence become increasingly strong with increasing environmental density and eventually dominate over the effects of shocks. In high-density regions, shock effects are negligible compared to turbulence. As for the missing baryons, a very small fraction originates from those expelled by feedback effects within galaxies (Connor et al., 2025). However, the vast majority are baryons that remain stranded outside galaxy clusters. Due to the hindrance of dynamical effects, they have never had the opportunity to accrete onto these structures. These dynamical effects result in high baryon fractions in low- and intermediate-density regions, but low baryon fractions in high-density regions. Furthermore, the cumulative mass fraction reveals that both the WHIM and warm gas are more extensively distributed in space.
  2. 2.
    For the thermodynamic aspect of turbulence, turbulent kinetic energy cascades from large-scale to small-scale eddies, dissipates into thermal energy, and heats the cosmic gas. Turbulence has significantly greater thermal effects than shocks, especially in regions with higher environmental dark matter densities. In low-density regions, turbulence’s thermal effects are about twice as large as those of shocks, and this ratio increases dramatically with density. This suggests that in virialized structures, turbulence dominates gas heating, making shock heating negligible. This finding calls into question the traditional view that shocks are the primary sources of heating in the ICM and IGM.
  3. 3.
@@
    It can be seen, from $z=1$ to $z=0$, that both the mean thermal energy density ($\overline{\varepsilon}_{\rm th}$) and the mean turbulent energy density ($\overline{\varepsilon}_{\rm turb}$) remain nearly constant, and the shapes and amplitudes of the environment-dependent wavelet energy spectra are essentially unchanged. These findings indicate that the injection and dissipation of turbulent energy have essentially reached a balance, and that the cosmic fluid is in a steady state within this redshift range.
  4. 4.
@@
    Due to turbulent heating, we see that as redshift decreases, an increasing amount of warm gas is heated and converted into the WHIM, and some even into hot gas. Consequently, the mass fraction of warm gas gradually decreases, while that of the WHIM increases. The WHIM is present in nearly all regions, with the majority concentrated in intermediate-density regions.
There are a large number of problems relevant to the study of the formation and evolution of galaxies and galaxy clusters. Among these, the study of the heating mechanisms in the ICM and IGM is particularly important. The absence of a sufficient heating mechanism leads to a phenomenon called cooling flow within the galaxy clusters and, in turn, to an excess of star birth in the galaxy, which is known as the overcooling problem (Voit, 2005). The main proposals to overcome this overcooling problem are heating mechanisms, such as SN feedback or AGN activities. In the popular semi-analytical models of galaxy formation, the heating mechanism is mainly based on SN and AGN feedback, and turbulent heating is not considered (see Guo et al., 2011; Henriques et al., 2015). However, as addressed in Silk and Nusser (2010), black holes cannot supply sufficient momentum in radiation to expel the gas via pressure alone.
To address the problem of overcooling, Weinberger et al. (2017) propose that a large fraction of the kinetic energy injected in the feedback mode is thermalized via shocks in the surrounding gas, which provides a distributed heating channel for black hole feedback, effectively suppressing star formation. Indeed, it is commonly believed that when density perturbations collapse, the associated gas in virialized halos is shock-heated to the virial temperature of the corresponding structure (Mo et al., 2010). However, based on our study, we find that turbulence contributes more than shocks to both the dynamical and thermodynamic effects of gas within galaxies.
Zhuravleva et al. (2014) studied the thermodynamic effects of turbulence in cosmic fluids. Through an investigation of the turbulent heating rate and radiative cooling rate in the ICM of the Perseus and Virgo clusters, the authors suggest that turbulent heating is sufficient to offset radiative cooling at each radius of the clusters, implying that turbulence is a powerful thermal source in the ICM. Their findings support the conclusions of our current work.
In summary, we find that, compared with turbulence in the cosmic fluid, shocks are unimportant in intermediate-density regions and even negligible in high-density regions, both dynamically and thermodynamically. This finding accounts for the origin of the WHIM in terms of both dynamics and thermodynamics, calls into question the traditional view of shock heating, and highlights the importance of turbulence in shaping the large-scale structure of the Universe, particularly in terms of the evolution of galaxies and galaxy clusters.
We have validated our main findings using data from TNG50-2, TNG100-1, WIGEON, and EAGLE simulations. In Appendices A and B we show that differences in the spatial resolution, box size, and sub-grid implementation of physical processes do not affect our main conclusions.
Acknowledgements.
@@
The authors thank the anonymous referee for helpful comments and suggestions. We acknowledge the use of the data from IllustrisTNG simulation for this work, and also acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).
Appendix A Comparison with TNG50-2 and TNG100-1
Figure 10: Same as Fig. 5 but for TNG50-2.
To assess the robustness of our findings against numerical resolution and box size, we employed the TNG50-2 and TNG100-1 simulations in addition to our primary TNG50-1 run. For TNG50-2, we utilized a regular $N^{3}_{g}$ grid with $N_{g}=1024$, yielding spatial and mass resolutions of $34.2h^{-1}{\rm kpc}$ and $3.4\times 10^{6}h^{-1}M_{\odot}$, respectively. For TNG100-1, we adopted the same grid configuration as in TNG50-1, which resulted in corresponding resolutions of $48.8h^{-1}{\rm kpc}$ and $1.0\times 10^{7}h^{-1}M_{\odot}$.
Figure 11: Same as Fig. 5 but for TNG100-1.
The comparison of Figs. 5 and 10 reveals only minor differences between TNG50-1 and TNG50-2, which indicates that our findings are essentially unaffected by the simulation resolution. The comparison of Figs. 10 and 11 shows that turbulence contributes nearly equally to both dynamical and thermodynamic effects in TNG50-2 and TNG100-1, unaffected by spatial resolution. However, there exists a significant difference in shock contribution between the two simulations. This discrepancy arises because TNG100-1 employs a larger simulation box size, resulting in a lower spatial resolution that is inadequate for properly resolving shock structures, which are thin spatial regions.
Appendix B Comparison with other simulations
We used WIGEON and EAGLE simulations to verify how different sub-grid physics affect the main results of our work.
WIGEON cosmological simulation is based on a five-order accuracy WENO scheme for the hydro-solver to compute the baryonic quantities such as temperature, density, and velocity, and incorporated with the standard particle-mesh method for the gravity calculation of dark matter particles (Feng et al. 2004; Zhu et al. 2013). Due to the five-order accuracy of its hydro-solver, this code is much effective to capture shockwave and turbulence structures of IGM. A uniform UV background is switched on at $z=11$ to mimic the reionization. The radiative cooling and heating processes are modeled with a primordial composition, following the method in Theuns et al. (1998). WIGEON does not include any sub-grid physics such as star formation, SN and AGN feedback into the simulations. By excluding the influence of these feedback processes, it is expected that the effects of turbulence can be exhibited to the most extent. The WIGEON simulations are evolved from the initial time to the present day in periodic cubical boxes with side length $50\,{h^{-1}{\rm Mpc}}$. With an equal number of grid cells and dark matter particles of $1024^{3}$, the space and mass resolutions are determined to be $48.8h^{-1}{\rm kpc}$ and $8.3\times 10^{7}{{\rm M}_{\odot}}$, respectively.
Figure 12: Same as Fig. 5 but for the WIGEON simulation.
The EAGLE project (Schaye et al. 2015) consists of a suite of cosmological, hydrodynamical simulations of a standard $\Lambda$ cold dark matter universe, with the cosmological parameters taken from the most recent Planck results. EAGLE runs with a modified version of the $N$-Body Tree-PM smoothed particle hydrodynamics code GADGET-3, and includes the sub-grid models such as radiative cooling, star formation, stellar mass-loss and metal enrichment, energy feedback from star formation, gas accretion on to, and mergers of, supermassive black holes, and AGN feedback. In this study we used the EAGLE data of the $100{\rm Mpc}$ run, which has the same particle numbers for both dark matter and baryonic matter ($N=1504^{3}$). As with TNG50-1, we also assigned these particles to a regular $1536^{3}$ grid, which yielded a spatial resolution of $44.1h^{-1}{\rm kpc}$ and a mass resolutions of $7.3\times 10^{6}h^{-1}{{\rm M}_{\odot}}$.
The comparison of Figs. 5 and 12 shows that, for both dynamical and thermodynamic effects, the contribution of shocks in the WIGEON simulation is significantly smaller than that in TNG50. Note that in the WIGEON simulation, there is no feedback from sub-grid physics, so these dynamical and thermodynamic effects should be entirely attributed to the contribution of turbulence, whereas the contribution from shocks can be completely neglected.
Figure 13: Same as Fig. 5 but for the EAGLE simulation.
For the EAGLE simulation, from Fig. 13, we observe that in low-density regions, the dynamical and thermodynamic effects of shocks are significantly smaller than those of turbulence. However, as the environmental density increases, both the dynamical and thermodynamic effects become progressively stronger compared to those in the TNG simulation. Nevertheless, the contribution of shocks remains smaller than that of turbulence, except for the reversal where the dynamical effects of shocks surpass those of turbulence at $\Delta_{{\rm dm}}>5\times 10^{3}$.
The EAGLE project employs the smoothed-particle hydrodynamics code GADGET-3 to perform the simulations. It is well established that smoothed-particle hydrodynamics, in contrast to Eulerian mesh-based schemes, is intrinsically less efficient at resolving turbulence and shocks. Consequently, EAGLE has to compensate for these deficiencies in the dynamical and thermodynamic effects of turbulence by enhancing feedback from sub-grid physics. These enhanced feedback mechanisms trigger strong secondary shocks. Consequently, shock-induced dynamical and thermodynamic effects appear markedly stronger in high-density regions than in the TNG runs. As can also be seen from Fig.6 of Chisari et al. (2018) and Fig.4 of Wang and He (2024a), EAGLE exhibits greater suppression of the power spectrum on scales of $k>10\,{h{\rm Mpc}^{-1}}$ (i.e., $\lambda<0.6\,{h^{-1}{\rm Mpc}}$), and lower baryon fraction in high-density regions than TNG. These indicate that the feedback effects of the sub-grid physics implemented in EAGLE should be stronger than those in TNG, which is consistent with our interpretation.
However, we note that EAGLE generally has weaker AGN feedback but stronger stellar feedback than TNG, as demonstrated by comparative studies of these two models (e.g., Davies et al. 2020; Oppenheimer et al. 2020; Habouzit et al. 2021; Ayromlou et al. 2023). Consequently, the strong feedback in EAGLE on small scales or in high-density regions should be primarily attributed to stellar feedback rather than AGN feedback.